Larry had a total of $500 he spent 20% of his money on a book
.how much did the book cost

Answer by ankor@dixie-net.com(22740)

Larry had a total of $500 he spent 20% of his money on a book
.how much did the book cost
:
Do this in your head, 20/100 = 1/5. 1/5 of 500 = $100